Analysis and Interpretation of Profitability Balance sheets and income statements for Costco Wholesale Corportation follow Consolidated Statements of Earnings September 1, September 2, August 28 For Fiscal Years Ended ($ millions) 2013 2012 2011 Revenue 97,062 $87,048 1,867 Net Sales $102,870 2,286 105,156 Membership fees 2,075 Total revenue 99,137 88,915 Operating expenses Merchandise costs Selling, general and administrative Preopening expenses 91,948 10,104 51 3,053 86,823 9,518 37 2,759 77,739 8,691 46 2,439 Operating Income Other income (expense) (99) 97 3,051 990 2,061 (95) 103 2,767 1,000 Interest expense 60 2,383 841 Interest income and other, net Income before income taxes Provision for income taxes Net income including noncontrolling interests Net income attributable to noncontrolling interests 1,767 (58) 1,709 1,542 (80) 1,462 (22) Net income attributable to Costco 2,039

Consolidated Balance Sheets September September (S millions, except par value and share data) 1, 2013 2, 2012 Assets Current assets Cash and cash equivalents Short-term investments Receivables, net Merchandise inventories Deferred income taxes and other current assets 4,644 1,480 1,201 7,894 621 15,840 3,528 1,326 1,026 7,096 550 13,526 Total current assets Property and equipment 4,409 11,556 4,472 585 21,022 (7,141) 13,881 562 30,283 4,032 10,879 4,261 374 19,546 (6,585) 12,961 653 27,140 Land Buildings and improvements Equipment and fixtures Construction in progress Less accumulated depreciation and amortization Net property and equipment Other assets Total assets Liabilities and equity Current liabilities Accounts payable Accrued salaries and benefits Accrued member rewards Accrued sales and other taxes Deferred membership fees Other current liabilities 7,872 7,303 2,037 1,832 710 382 1,167 1,089 13,257 397 1,101 966 12,260 Total current liabilities

13,257 4,998 1,016 19,271 Total current liabilities 12,260 1,381 981 14,622 Long-term debt, excluding current portion Deferred income taxes and other liabilities Total liabilities Equity Preferred stock, $0.005 par value 100,000,000 shares authorized; no shares issued and outstanding Common stock, $0.005 par value 900,000,000 shares authorized; 436,839,000 and 432,350,000 shares issued and outstanding Additional paid-in-capital Accumulated other comprehensive (loss) income Retained earnings Total Costco stockholders equity 4,670 (122) 6,283 10,833 179 11,012 30,283 4,369 156 7,834 12,361 157 12,518 $27,140 Noncontrolling interests Total equity Total liabilities and equity

HINT: For Sales use Total revenues for your computations, when applicable. (a) Compute net operating profit after tax (NOPAT) for 2013. Assume that the combined federal and state statutory tax rate is 37%.(Round to the nearest whole number.) 2013 NOPAT 2,019 x (S millions) (b) Compute net operating assets (NOA) for 2013 and 2012. 2013 NOA10,454x($ millions) 2012 NOA: 11,936 ($ millions) (c) Compute Costcos RNOA, net operating profit margin (NOPM) and net operating asset turnover (NOAT) for 2013. (Do not round until final answer. Round two decimal places. Do not use NOPM x NOAT to calculate RNOA) 2013 RNOA 18.04 2013 NOPM-1.96 2013 NOAT9.19 (d) Compute net nonoperating obligations (NNO) for 2013 and 2012. 2013 NNO-0 2012 NNO-0 x$millions) x$millions) (e) Compute return on equity (ROE) for 2013. (Do not round until final answer. Round answer two decimal places) 2013 ROE-18.64 (f) Infer the nonoperating return component of ROE for 2013. (Use answers from above to calculate. Round two decimal places.) 0.88 (g) What does the relation between ROE and RNOA suggest about Costcos use of equity capital? ROE RNOA implies that Costcos equity has grown faster than its NOA ROE> RNOA implies that Costco has taken on too much financial leverage.x ROE> RNOA implies that Costco is able to borrow money to fund operating assets that yield a return greater than its cost of debt ROE RNOA implies that Costco is holding significant levels of nonoperating assets that are not earning as big of a return as its operating assets.

Please help me answer these correctly with solutions. Thank you!
Answer #1

a.. NOPAT:

Operating income 2013 = 3053

Tax = 37%

Therefore, NOPAT = Operating income * (1-Tax rate) = 3053* (1-0.37) = 1923.39 milliions

b. Net Operating Assets = Total Assets - All liabilities - All financial assets + All financial liabilities

For 2013:

Total assets = $30283, All liabilities = $19271, Financial assets = cash and cash equivalent + short term investments = 4644 + 1480 = 6124, Financial liabilities = Long term debt, excluding current portion = 4998

Therefore NOA for 2013 = $11366 millions

For 2012, NOA = $9045 millions

c. 2013 RNOA = NOPAT / NOA = 1923.39 / 11366 = 16.9%

2013 NOPM = NOPAT / Total revenues = 1923.39 / 105156 = 1.8%

2013 NOAT = Total revenue / average NOA = 105156 / ((11366+9045)/2) = 105156 / 10205.5 = 10.30

d. Net nonoperating obligations = Non operating liabilities - Non operating assets

For 2013, Non operating liabilities = Financial liabilities = Long term debt, excluding current portion = 4998

Non operating assets = cash and cash equivalent + short term investments = 6124

NNO = 4998 - 6124 = - $1126 millions

For 2012, NNO = 1381 - (3528+1326) = $3473 millions

e. Return on equity = Net income / shareholder's equity average

For Costco, Net income 2013 excluding non controlling interest = 2039

Total Costco shareholder's equity = 10833

2013 ROE = 2039 / 10833 = 18.88%

Including non controlling interest = 2013 ROE = Total income including non controlling interest / Total equity = 2061 / 11012 = 18.71%

f. Non operating 2013 ROE = (Total income including non controlling interest - NOPAT) / Total equity = (2061 - 1923.39) / 11012 = 1.25%

g. ROE > RNOA implies Costco is able to borrow money to fund operating assets that yield a return greater than its cost of debt